2x+y=25


AND
x-y=2

Answer by rfadrogane(214)

2x+y=25 AND x-y=2
Sol'n:
add the 2 equation:
3x = 27
x = 9--answer
from x-y=2,
y = x - 2
= 9-2
y = 7---answer
^_^ even though you did not put your question, i think this what you me to do